OpenStudy (anonymous):

-11+2b≤5 (Show Work)

OpenStudy (anonymous):

add 11 and divide by 2

OpenStudy (anonymous):

\[-11 + 2b \le 5\]2b add 11 to both sides

OpenStudy (anonymous):

we are now left with \[2b \le 16\]

OpenStudy (anonymous):

Now divide both sides by 2

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

we now have \[b \le 8\]

OpenStudy (anonymous):

Sorry for the slow reply. i have a really poor intrnet connection

OpenStudy (anonymous):

Sorry for the slow reply. i have a really poor internet connection

OpenStudy (anonymous):

sorry. I know that sucks.